On a graph paper , take 2 cm to represent one unit on both axes, draw the lines:
x+3=0,y-2=0,2x+3y=12.
write down the coordinates of the vertices of the triangle

Answer by Alan3354(69443)

On a graph paper , take 2 cm to represent one unit on both axes, draw the lines:
x+3=0,y-2=0,2x+3y=12.
write down the coordinates of the vertices of the triangle
======================
You can use any scale on the graph paper, just graph the 3 lines.
-----
You can solve for the intersections algrebraically. Take them 2 at a time.